OpenStudy (anonymous):

integral (x^2-1)/(x^3+1)

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ x ^{2}-1 }{ x ^{3}+1 }\]

OpenStudy (anonymous):

no (x^2-1)/(x^3+1)

OpenStudy (amoodarya):

I think "Baseballa101" were wrong !

OpenStudy (anonymous):

x^2-1=(x-1)(x+1) and x^3+1=(x+1)(x^2-x+1) use the above 2 equations and try to solve. this might help. :)

OpenStudy (amoodarya):